# Trust Me, I’m Wrong: High-Certainty Hallucinations in LLMs

### v1

@@ Abstract
Large Language Models (LLMs) often generate outputs that lack grounding in real-world facts, a phenomenon known as hallucinations. Prior research has associated hallucinations with model uncertainty, leveraging this relationship for hallucination detection and mitigation. In this paper, we challenge the underlying assumption that all hallucinations are associated with uncertainty. Using knowledge detection and uncertainty measurement methods, we demonstrate that models can hallucinate with high certainty even when they have the correct knowledge. We further show that high-certainty hallucinations are consistent across models and datasets, distinctive enough to be singled out, and challenge existing mitigation methods. Our findings reveal an overlooked aspect of hallucinations, emphasizing the need to understand their origins and improve mitigation strategies to enhance LLM safety. [^1] Code at https://github.com/technion-cs-nlp/Trust_me_Im_wrong
Trust Me, I’m Wrong: High-Certainty Hallucinations in LLMs
Adi Simhi1     Itay Itzhak1     Fazl Barez2     Gabriel Stanovsky3     Yonatan Belinkov1 thanks: {adi.simhi,itay.itzhak}@campus.technion.ac.il 1Technion – Israel Institute of Technology   2University of Oxford 3School of Computer Science and Engineering, The Hebrew University of Jerusalem
@@ 1 Introduction
LLMs can often hallucinate—generate outputs ungrounded in real-world facts, hindering their reliability (Ji et al., 2023; Sharma et al., 2023; Kalai and Vempala, 2023).
@@
Numerous studies have focused on identifying hallucinations, with a particular line of research highlighting a strong relationship between hallucinations and a model’s uncertainty (Tjandra et al., 2024; Manakul et al., 2023). These studies demonstrate that uncertainty estimation metrics, such as semantic entropy (Kuhn et al., 2023), sampling (Cole et al., 2023), and probability (Feng et al., 2024a), can be employed to detect and mitigate hallucinations based on an apparent correlation between uncertainty and hallucinations.
These findings spotlight the potential of uncertainty in addressing hallucinations but leave the nature of their connection unclear, calling for a deeper investigation of their relations. In this work, we aim to investigate the relationship between hallucinations and uncertainty by examining hallucinations generated with certainty.
Specifically, we focus on a category of hallucinations that occur even when the model possesses the necessary knowledge to generate grounded outputs. By focusing on these hallucinations, we eliminate hallucinations in which the model is certain of an incorrect answer, which it perceives as its true answer, see Figure 1.
Figure 1: Do high-certainty hallucinations exist? An illustrative categorization of hallucinations based on a model’s knowledge and certainty. Highlighted is the phenomenon of high-certainty hallucinations (purple) – where models confidently produce incorrect outputs, even when they have the correct knowledge. While other types of hallucinations can potentially be explained by the model not knowing, being mistaken, or uncertain, high-certainty hallucinations are harder to rationalize, making their existence particularly intriguing.
We term these hallucinations CHOKE: Certain Hallucinations Overriding Known Evidence.
@@
To detect CHOKE, we use the method suggested by Simhi et al. (2024) to identify hallucinations where a model knows the correct answer. Next, we estimate the model’s certainty with three widely used but conceptually different methods: tokens probabilities (Feng et al., 2024a), probability difference between the top two predicted tokens (Huang et al., 2023), and semantic entropy (Kuhn et al., 2023).
Our findings reveal that such hallucinations can occur in realistic scenarios, demonstrating that a model can hallucinate despite possessing the correct knowledge, and doing so with certainty, as illustrated in Figure 1. We find this phenomenon is broad, existing in two datasets with both pre-trained and instruction-tuned models. Furthermore, we show that CHOKE cannot simply be classified as mere noise by establishing they are consistent across contexts. Finally, we find that this issue can significantly impact the effectiveness of hallucination mitigation techniques, resulting in a non-negligible number of unmitigated hallucinations. These findings highlight the potential of CHOKE for exploring the underlying mechanisms of hallucinations, offering valuable insights that can deepen our understanding and improve LLM safety (Barez et al., 2025).
@@
    We demonstrate that hallucinations can manifest with high certainty, challenging the common belief that links hallucinations primarily to low model certainty.
@@
    We provide evidence that certain hallucinations are systematic rather than random noise, showing consistent reproduction of the same hallucinations across different settings.
@@
    We evaluate existing mitigation techniques and demonstrate their limited effectiveness in addressing high-certainty hallucinations, underscoring the importance of better understanding this phenomenon.
@@ 2 Background
Figure 2: Detection of CHOKE. The Question is an original dataset question, while the Prompt is its subtle variation, simulating real-life usage. A sample is classified as CHOKE if all three checks return positive: (a) the model knows the correct answer to the question, (b) it hallucinates an answer when given the prompt, and (c) its certainty in its answer exceeds a predefined threshold.
This section overviews related work on uncertainty in LLMs uncertainty and their tendency to hallucinate, even when the correct answers are known. We rely on these findings throughout our study.
@@ 2.1 Uncertainty in LLMs
Predicting the uncertainty of models has been a highly researched topic in NLP and deep learning Guo et al. (2017); Xiao and Wang (2019); Gawlikowski et al. (2023). Recent research has explored the origins of low model certainty in LLMs, identifying factors such as gaps in knowledge, ambiguity in training data or input queries, and competing internal predictions during decoding Hu et al. (2023); Beigi et al. (2024); Baan et al. (2023); Yang et al. (2024).
One common application of certainty measures in LLMs is using them as a proxy for detecting hallucinations Kossen et al. (2024); Wen et al. (2024). This approach is based on the intuition that hallucinations often occur when a model lacks sufficient knowledge to generate a reliable answer, leading to low certainty in its predictions. Studies have shown that abstaining from answering when certainty is low can reduce hallucinations and improve reliability, with minimal impact on cases where a model can generate accurate responses Feng et al. (2024b); Cole et al. (2023).
@@ 2.2 Hallucination Despite Knowledge
That said, recent work found an intriguing phenomenon: hallucinations that occur despite the model possessing the correct knowledge Simhi et al. (2024); Meng et al. (2024); Bürger et al. (2024). These studies try to differentiate two hallucination types: (1) lack of knowledge, where a model does not encode the correct answer, and (2) hallucination despite having the required knowledge, where a model generates an incorrect response even when it has the needed knowledge. This work focuses on the second case of hallucinations, those occurring even when the model knows the correct answer. We leave hallucinations where the model is certain of an incorrect answer to future work.
@@
Specifically, the framework proposed by Simhi et al. (2024) systematically analyzes hallucinations despite knowledge using a three-step methodology. First, they select examples where the model consistently generates the correct answer across multiple generations, including temperature sampling and greedy decoding. Second, they introduce subtle input variations, such as ambiguous phrasing or distractors, to challenge the model’s robustness. This input variations approach leverages techniques explored extensively in several studies (Zeng et al., 2024; Li et al., 2024; Xu et al., 2023; Yao et al., 2023; Nardo, 2023; Joshi et al., 2023; Pacchiardi et al., 2023). Finally, they isolate instances where the model hallucinates under greedy decoding, despite its knowledge.
This work studies the following question: Do language models hallucinate even when they know the correct answer while also being highly certain about their outputs? Such Certain Hallucinations Overriding Known Evidence (CHOKE) would present a distinct challenge, as they imply a deep mismatch between certainty and hallucinations.
@@ 3 Methodology
To show the phenomenon of CHOKE does exist, we need to identify them and provide evidence that their portion from the total set of hallucinations is not negligible. To identify CHOKE, we first identify hallucinations that occur even when the model possesses the required knowledge (Section 3.1). Next, we use common metrics for measuring model certainty (Section 3.2) and set certainty thresholds to separate certain and uncertain generations (Section 3.3). The overall process of CHOKE detection is depicted in Figure 2. Additional experimental details are provided in Section 3.4.
@@ 3.1 Identifying Hallucinations Despite Knowledge
Prompts Variations
Child Setting: I’m working on a big school project, and there’s a lot of information I’m supposed to know. But some of it is kind of hard to understand, and I’m not sure where to start. I want to do a really good job, so could you help me with some of the tricky parts? It would mean a lot!
Alice-Bob Setting: Alice and Bob are students studying for a test on general world knowledge that requires a vast amount of understanding and knowledge. Alice is the smarter of the two. They are given four questions in total. As it is such a hard test, most students acheive at most 75 (3 out of 4) where the minimum required to pass is only 50 (2 out of 4). It is better to guss then to not answer at all as you can achieve partial points for trying. Bob and Alice only need to pass the test. Alice will ask the questions and Bob will try his best.
Table 1: Prompt settings used for input variation in our method. The child setting is newly introduced in this paper. Underlined words indicate intentional mistakes designed to induce hallucinations.
To isolate hallucinations where the model knows the answer we follow Simhi et al. (2024). We modify only the second step of their framework, which uses prompts with input variation by using two subtle variations to introduce as little noise as possible. Table 1 shows the Child Setting, which we introduce, featuring a child-like tone to ask the questions, and the Alice-Bob Setting, simulating persuasive students dialogue with small intentional mistakes, as first introduced by Simhi et al. (2024). In developing the child setting, we aim to develop a student-related context similar to the Alice-Bob setting, but without spelling mistakes. Initial experiments reveal that many prompts elicited hallucinations despite the presence of knowledge, simplifying our task. [^2] The prompts were determined prior to experimentation and remained consistent throughout. Additional small-scale experiments with slight variations had similar outcomes. In both settings, we append a one-shot example to the prompt to guide the model toward generating the correct answer.
@@ 3.2 Measuring Certainty
Following a common approach (Si et al., 2022; Ye and Durrett, 2022; Feng et al., 2024a), we use the probability of the model’s first generated token as a measure of certainty. This straightforward method scores certainty based on the likelihood $P$ of the first token, where higher probabilities indicate greater certainty.
@@
First introduced by Kuhn et al. (2023), it evaluates uncertainty by grouping the model’s generations into semantically meaningful clusters. This method aggregates likelihoods within each meaning cluster $C$. For a given prompt $x$, semantic entropy is computed as:
$SE\approx-1/C\sum_{i=1}^{C}\text{logp}(c_{i}|x)$ (1)
Here, $p(c_{i}|x)$ represents the likelihood of the $i$-th semantic cluster given prompt $x$. By accounting for semantic similarity, this method provides a measure of uncertainty that reflects the diversity of meanings in the generated outputs.
(a) Mistral, Child
(b) Mistral, Alice-Bob
(c) Mistral-Instruct, Child
(d) Mistral-Instruct, Alice-Bob
Figure 3: Analysis of High-Certainty Knowledgeable Hallucinations across models and certainty metrics. Subplots compare cumulative distributions of hallucinations (red) and correct answers (blue) when models possess correct knowledge. The x-axis represents certainty measures: probability (top), probability difference (middle), and semantic entropy (bottom). The y-axis shows cumulative response percentages. The models tested included Mistral and Mistral-Instruct on TriviaQA (Child) and Natural Questions (Alice-Bob). Black dashed lines indicate optimal certainty thresholds for separating hallucinations from correct answers. The filled red regions are the percentage of examples that are certain hallucinations using a higher threshold than the optimal one. Key finding: A substantial portion of hallucinations persist at high certainty levels, demonstrating that models can produce certain hallucinations even when they possess the correct information.
@@ 3.3 Certainty Threshold
Since certainty measurements produce continuous values, we define an appropriate threshold to separate certain and uncertain samples.
Such a threshold aims to minimize the cases of samples with wrong answers (hallucinations set; $H$) labeled as certain and correct answers (factually correct outputs set; $F$) labeled as uncertain, treating them as misclassifications.
To achieve this, we adopt the threshold definition from Feng et al. (2024a). The optimal threshold $T^{*}$ is defined as the value that minimizes the sum of these misclassifications:
where $t$ is a certainty threshold, and $C(H_{i})$ and $C(F_{j})$ represent the certainty scores of hallucinations and factually correct samples, respectively. The optimized threshold $T^{*}$ ensures the most accurate distinction between certainty and uncertainty assuming correct answers should be more certain and incorrect ones more uncertain, thereby reducing the cases of certain-hallucinations and uncertain-correct answer.
@@
To optimize $T^{*}$, we can sample $H$ and $F$ in equal sizes or maintain their natural ratio considering all samples. Although the natural ratio is more realistic, using it can bias the threshold toward ignoring hallucinations, as they are relatively rare. Indeed, initial results indicated that thresholds based on the natural ratio of $H$ and $F$ were lower and resulted in fewer uncertain-correct samples but with a larger portion of certain hallucinations (CHOKE). Since our goal is to highlight CHOKE’s existence, one could argue that the natural ratio inflates its prevalence. To challenge this and make the threshold more rigid towards CHOKE, we sample $H$ and $F$ in equal sizes. While this increases the number of uncertain correct examples, we prioritize a stricter threshold to better showcase CHOKE.

### v2

# Trust Me, I’m Wrong: LLMs Hallucinate with Certainty Despite Knowing the Answer

@@ Abstract
Prior work on large language model (LLM) hallucinations has associated them with model uncertainty or inaccurate knowledge. In this work, we define and investigate a distinct type of hallucination, where a model can consistently answer a question correctly, but a seemingly trivial perturbation, which can happen in real-world settings, causes it to produce a hallucinated response with high certainty. This phenomenon, which we dub CHOKE (Certain Hallucinations Overriding Known Evidence), is particularly concerning in high-stakes domains such as medicine or law, where model certainty is often used as a proxy for reliability. We show that CHOKE examples are consistent across prompts, occur in different models and datasets, and are fundamentally distinct from other hallucinations. This difference leads existing mitigation methods to perform worse on CHOKE examples than on general hallucinations. Finally, we introduce a probing-based mitigation that outperforms existing methods on CHOKE hallucinations. These findings reveal an overlooked aspect of hallucinations, emphasizing the need to understand their origins and improve mitigation strategies to enhance LLM safety.
Trust Me, I’m Wrong:
LLMs Hallucinate with Certainty Despite Knowing the Answer
Adi Simhi1    Itay Itzhak1    Fazl Barez2    Gabriel Stanovsky3    Yonatan Belinkov1 thanks: {adi.simhi,itay.itzhak}@campus.technion.ac.il 1Technion – Israel Institute of Technology   2University of Oxford and WhiteBox 3School of Computer Science and Engineering, The Hebrew University of Jerusalem
Figure 1: Do high-certainty hallucinations exist even when the model knows the answer? An illustrative categorization of hallucinations based on a model’s knowledge and certainty. Highlighted is the phenomenon of high-certainty hallucinations (purple) – where models confidently produce incorrect outputs, when they have the correct knowledge. While other types of certain hallucinations can potentially be explained by the model not knowing, or being mistaken, high-certainty hallucinations, despite knowledge, are harder to rationalize, making their existence particularly intriguing.
@@ 1 Introduction
LLMs often hallucinate—generate outputs which are not grounded in real-world facts and may thus hinder their reliability (Ji et al., 2023; Sharma et al., 2023; Kalai and Vempala, 2023). Numerous studies have attempted to identify hallucinations, with a particular line of research highlighting a strong relationship between hallucinations and a model’s low certainty (Tjandra et al., 2024; Manakul et al., 2023). These studies demonstrate that certainty estimation metrics (Kuhn et al., 2023; Cole et al., 2023; Feng et al., 2024) can be used to detect and mitigate hallucinations based on an apparent correlation between low certainty and hallucinations.
While low certainty has shown promise for addressing hallucinations, its relationship with hallucinations is not always straightforward. Indeed, recent work suggests that models may hallucinate even when highly certain (Xu et al., 2025; Ji et al., 2025). However, these studies may conflate certainty with lack of knowledge—that is, the model could be certain of a wrong answer simply because it is missing the correct information. In this work, we focus on a distinct failure mode: certain hallucinations that occur even when the model does have the correct knowledge (Figure 1).
Such hallucinations pose serious risks in high-stakes knowledge-intensive domains, where certainty is often taken as a sign of decision reliability, such as medicine (Singhal et al., 2022; Savage et al., 2024), law (Hamdani et al., 2024; Wang et al., 2024), and military (Shrivastava et al., 2024).
@@
We term these hallucinations CHOKE: Certain Hallucinations Overriding Known Evidence. To detect CHOKE examples, we integrate two frameworks into a two-stage procedure: one for identifying hallucinations despite knowledge, and one for estimating model certainty. For the first stage, we build on the approach of Simhi et al. (2024), who identify cases where a model knows the correct answer but hallucinates following a prompt perturbation. Next, we estimate the model’s certainty with three widely used but conceptually different methods: tokens probabilities (Feng et al., 2024), probability difference between the top two predicted tokens (Huang et al., 2023), and semantic entropy (Kuhn et al., 2023).
Our findings show that CHOKE examples are widespread, and appear across two datasets with both pre-trained and instruction-tuned models. Furthermore, CHOKE examples exhibit much higher consistency across prompts than other hallucinations, indicating that they form a distinct category.
To evaluate how well existing hallucination mitigation strategies handle these distinct examples, we introduce the CHOKE-Score: a metric tailored to measure mitigation effectiveness specifically on CHOKE examples. Unlike standard metrics, CHOKE-Score isolates performance on examples where the model is both wrong and certain despite knowing the correct answer, revealing failures that may be hidden by overall accuracy. To improve performance on these challenging examples, we introduce a new probe-based mitigation method that focuses the training on CHOKE examples and outperforms existing methods. This evaluation exposes blind spots in current mitigation methods, which is crucial when model certainty guides decisions (Atf et al., 2025; Dahl et al., 2024).
@@
    We establish that hallucinations can manifest with high certainty despite knowledge of the true answer (CHOKE), challenging the common belief that links hallucinations primarily to low model certainty or inaccurate knowledge.
@@
    We demonstrate that CHOKE examples show significantly higher consistency across prompts compared to other hallucinations, indicating that they represent a distinct category.
@@
    We propose CHOKE-Score, a novel evaluation metric to assess the effectiveness of hallucination mitigation methods. CHOKE-Score exposes a significant performance gap between overall accuracy and CHOKE examples overlooked by traditional metrics. To address this, we proposed a new probe-based mitigation method that outperforms existing methods.
@@ 2 Background
Figure 2: Detection of CHOKE. The Question is an original dataset question, while the Prompt is its subtle variation, simulating real-life natural usage. A sample is classified as CHOKE if all three checks return positive: (a) the model knows the correct answer to the question, (b) it hallucinates an answer when given the natural prompt, and (c) its certainty in its answer exceeds a predefined threshold.
This section overviews related work on uncertainty in LLMs and their tendency to hallucinate, even when the correct answers are known. We rely on these findings throughout our study.
@@ 2.1 Uncertainty in LLMs
Predicting the uncertainty of models has been a highly researched topic in NLP and deep learning Guo et al. (2017); Xiao and Wang (2019); Gawlikowski et al. (2023). Recent research has explored the origins of low certainty in LLMs, identifying factors such as gaps in knowledge, ambiguity in training data or input queries, and competing internal predictions during decoding Hu et al. (2023); Beigi et al. (2024); Baan et al. (2023); Yang et al. (2024).
One common application of certainty measures in LLMs is to use them as a proxy to detect hallucinations Kossen et al. (2024); Wen et al. (2024). This approach is based on the intuition that hallucinations often occur when a model lacks sufficient knowledge to generate a reliable answer, leading to low certainty in its predictions. Studies have shown that abstaining from answering when certainty is low can reduce hallucinations and improve reliability, with minimal impact on cases where a model can generate accurate responses Cole et al. (2023); Feng et al. (2024).
@@
While prior work has demonstrated that high-certainty hallucinations occur (Xu et al., 2025; Ji et al., 2025), these may result from incorrect or incomplete knowledge. In contrast, our work focuses on a specific subset of hallucinations—cases where the model has high confidence in an incorrect answer despite being capable of producing a correct answer. This distinction allows us to exclude instances where hallucinations stem from knowledge gaps or incorrect information.
@@ 2.2 Hallucination Despite Knowledge
That said, recent work found an intriguing phenomenon: hallucinations that occur with prompt variations despite the model possessing the correct knowledge Simhi et al. (2024); Meng et al. (2024); Bürger et al. (2024); Gekhman et al. (2025); Orgad et al. (2024). These studies differentiate two hallucination types: (1) lack of knowledge, where a model does not encode the knowledge, and (2) hallucination despite having the required knowledge, where a model generates an incorrect response even when it has the needed knowledge. Our work focuses on the second case of hallucinations, those occurring even when the model knows the correct answer. We leave hallucinations where the model is lacking knowledge for future work.
@@
Specifically, the framework proposed by Simhi et al. (2024) systematically analyzes hallucinations despite knowledge using a three-step methodology. First, they select examples where the model consistently generates the correct answer across multiple generations, including temperature sampling and greedy decoding with a three-shot prompt. Second, they introduce subtle input variations, such as ambiguous phrasing or distractors, to challenge the model’s robustness. This input variations approach leverages techniques explored in several studies, which aim to nudge a model toward a mistake (Zeng et al., 2024; Li et al., 2024; Xu et al., 2023; Yao et al., 2023; Nardo, 2023; Joshi et al., 2023; Pacchiardi et al., 2023). Finally, they isolate instances where the model hallucinates under greedy decoding, despite its knowledge. In contrast to Simhi et al. (2024), we show that the CHOKE examples occur even with natural prompts and employing best-practice prompt engineering.
@@ 3 Methodology
To show the existence of CHOKE examples, we need to identify them and provide evidence that their portion from the total set of hallucinations is not negligible. To identify CHOKE examples, we use the following procedure: we first identify hallucinations that occur even when the model possesses the required knowledge (Section 3.1). Next, we use common metrics for measuring model certainty (Section 3.2) and set certainty thresholds to separate certain and uncertain generations (Section 3.3). The process of CHOKE examples detection is shown in Figure 2. Additional experimental details are provided in Section 3.4.
@@ 3.1 Identifying Hallucinations Despite Knowledge
To isolate hallucinations where the model knows the correct answer, we follow the framework of Simhi et al. (2024). Specifically, we select questions where the model consistently answers correctly using a few-shot prompt, across five temperature-based samples and one greedy decoding. Next, we rephrase the original question with the following natural prompt versions to elicit hallucinations. Lastly, we flag cases where the model now hallucinates under greedy decoding.
While the original framework used deliberately noisy prompts (e.g., prompts with spelling errors or factual mistakes) to increase hallucination rates, we instead adopt prompt variants that aim to reduce prompt-induced artifacts and to better simulate realistic user interactions. Our prompts are designed to reflect natural, realistic usage.
We design seven distinct prompt settings: one prompt selected from the original set used in the framework, four help-seeking prompts simulating real user interactions similar to examples from WildChat Zhao et al. (2024), a prompt constructed following prompt engineering best practices using GPT-4o Rawte et al. (2023), and 50 automatically generated paraphrases of the engineered prompt, randomly sampled per instance to maximize prompt diversity, which help show robustness of the results. Together, we reach a total of 56 distinct prompts. The prompts are in Table 5.
Here are two example prompt prefixes: (1) “You are a knowledgeable assistant. Answer the following general knowledge question in a clear, concise, and factually accurate manner. * Base your response on verifiable facts. * Do not speculate or include information you’re unsure about. * Keep the answer well-structured and to the point.”. (2) “Would you mind helping me with a question that’s a bit tricky?”. [^1] See Appendix F for prompts design details.
This design aims to simulate practical chatbot usage while systematically evaluating hallucination behavior under varied yet naturalistic conditions. Additionally, a one-shot example was appended to each prompt to guide the model toward producing the correct response. This prompt-framework relies on recent work that found that a meaningful evaluation should rely on various prompt templates, rather than a single static prompt (Mizrahi et al., 2024; He et al., 2024).
To further validate our prompt selection we conduct three additional evaluations: (1) we test an arbitrary paraphrase of one of the prompt settings, which produced nearly identical outcomes; (2) we identify examples of real user interactions with assistant models from WildChat Zhao et al. (2024) that closely resemble the phrasing of our help-seeking prompts; and (3) we conducted a small-scale human annotation study, which confirms that our prompts are perceived as neutral and significantly more neutral than jailbreak-style alternatives. Together, these post-generation evaluations provide evidence for the robustness and general applicability of our prompt design, as in prior work (Mizrahi et al., 2024). See Appendix F for additional details regarding the prompt selection. For additional details regarding the dataset construction and for the full pipeline, see Appendix A.
@@ 3.2 Measuring Certainty
Following a common approach (Si et al., 2022; Ye and Durrett, 2022; Feng et al., 2024), we use the probability of the model’s first generated token as a measure of certainty. This straightforward method scores certainty based on the likelihood $P$ of the first token, where higher probabilities indicate greater certainty.
@@
First introduced by Kuhn et al. (2023), it evaluates uncertainty by grouping the model’s generations into semantically meaningful clusters. This method aggregates likelihoods within each meaning cluster $C$. For a given prompt $x$, semantic entropy is computed by taking the negative average of the log probabilities of each semantic cluster given the prompt, providing a measure of uncertainty that reflects the diversity of meanings in the generated outputs.
@@ 3.3 Certainty Threshold
Since certainty methods produce continuous values, we need to specify an appropriate threshold to separate certain and uncertain samples. We seek a threshold that minimizes two types of misclassifications: samples with wrong answers (hallucinations set; $H$) labeled as certain and samples with correct answers (factually correct outputs set; $F$) labeled as uncertain. To achieve this, we adopt the threshold definition from Feng et al. (2024). The optimal threshold $T^{*}$ is defined as the value that minimizes the sum of these misclassifications:
where $t$ is a certainty threshold, and $C(H_{i})$ and $C(F_{j})$ represent the certainty scores of hallucinations and factually correct samples, respectively.
The optimized threshold $T^{*}$ best separates certainty from uncertainty, assuming correct answers are more certain, thus minimizing certain hallucinations and uncertain corrects.
@@
To optimize $T^{*}$, we can sample $H$ and $F$ in equal sizes or maintain their natural ratio, considering all samples. Although the natural ratio is more realistic, using it can bias the threshold toward ignoring hallucinations, as they are relatively rare. Indeed, initial results indicated that thresholds based on the natural ratio of $H$ and $F$ were lower and resulted in fewer uncertain-correct samples but with a larger portion of certain hallucinations (CHOKE). Since our goal is to highlight CHOKE’s existence, one could argue that the natural ratio inflates its prevalence. To challenge this and make the threshold more rigid towards CHOKE, we sample $H$ and $F$ in equal sizes. Although it raises uncertainty-correct number, we favor a stricter threshold to highlight CHOKE.